OpenStudy (anonymous):

a cone has circular base of radius 5cm and a hieght of 12cm. find the curved surface area

OpenStudy (wolf1728):

If radius = 5 and height = 12 then (using the Pythagorean Theorem) the slant height = square root (12^2+ 5^2) = square root (169) = 13 The formula for the lateral area (curved surface area) can be found here: http://www.1728.org/volcone.htm Lateral Area = (PI * r * slant height) Lateral Area = 3.14159 * radius * 13 Lateral Area = 3.14159 * 5 * 13 Lateral Area = 204.20

OpenStudy (dan815):